vector A of magnitude 20 units lies in the direction 45 degrees S of E, while vector B, of magnitude 30 units is in the direction 60 degrees W of N. Calculate the scaler product A.B


1
Expert's answer
2020-12-14T12:13:01-0500

"\vec{A}*\vec{B}=| \vec{A}|| \vec{B}|\cos\theta"

"\theta_1= \angle(W,S)+45^0\text{(S of E)}+60^0\text{(W of N)}="

"= 90^0+45^0+60^0= 195^0"

"\theta_1>180^0\text{ hence this is the outer angle}"

"\theta_1= 360^0-\theta= 165^0"

"\vec{A}*\vec{B}=| \vec{A}|| \vec{B}|\cos\theta=20*30*\cos{165^0}\approx-579.6"

Answer: -579.6 scaler product